4.1.A supplier shipped a lot of six parts to a company. The lot contained three defective parts. Suppose the customer decided to randomly select two parts and test them for defects. How large a sample space is the customer potentially working with? List the sample space. Using the sample space list, determine the probability that the customer will select a sample with exactly one defect.
Answer by stanbon(75887)

A supplier shipped a lot of six parts to a company. The lot contained three defective parts. Suppose the customer decided to randomly select two parts and test them for defects.
---------------------------------------
How large a sample space is the customer potentially working with?
# of elements in the sample space: 6C2 = 15
I will not list them.
---------------------------
Using the sample space list, determine the probability that the customer will select a sample with exactly one defect.
---
# of elements in the sample space: 6C2 = 15
I will not list them.
-----
P(2 defects) = (3/6)^2 = 1/4
P(no defects) = (1/2)^2 = 1/4
So, P(1 defect) = 1- [1/4 + 1/4] = 1/2
------------------------
Cheers,
Stan H.
==================